Circular No. 7273

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 T1 (McNAUGHT-HARTLEY) Robert H. McNaught, Research School of Astronomy and Astrophysics,

Siding Spring Observatory, reports his discovery of a comet on a plate taken by Malcolm Hartley with the 1.2-m U.K. Schmidt Telescope. The strongly condensed comet showed a 8′′ coma and a very faint 1′ tail in p.a. 320^{o}.

1999 UT  α_{2000}  δ_{2000}  m_{1}
Oct. 7.63654  2^{h}04^{m}29.^{s}37  −39^{o}08′08′′.0  15
7.68168  2 04 27.46  −39 08 31.1

Additional astrometry is published on MPEC 1999-T42. I. P. Griffin, Auckland Observatory, reports a condensed coma of diameter 7′′ on CCD exposures taken on Oct. 11.4 UT (0.5-m telescope). A. Becker and C. Stubbs, University of Washington; and J. Perez, Cerro Tololo Interamerican Observatory (CTIO) note a tail in p.a. 328^{o} on a CCD exposure taken with the CTIO 0.9-m telescope on Oct. 11.19.

Preliminary parabolic orbital elements from MPEC 1999-T42:

T = 2000 Dec. 9.634 TT  ω = 345.^{o}173
Ω = 181.735 }2000.0
q = 1.15380 AU  i = 79.094
1999TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Oct. 9 2^{h}03.^{m}48 −39^{o}20.^{′}5 4.608 5.329 132.^{o}1 8.^{o}0 15.6
14 1 59.62 −40 03.8 4.575 5.283 130.9  8.2 15.5
19 1 55.53 −40 43.1 4.548 5.236 129.3  8.5 15.5
24 1 51.27 −41 17.9 4.527 5.189 127.2  8.8 15.4
29 1 46.91 −41 47.9 4.511 5.142 124.7  9.1 15.4
Nov. 3 1 42.51 −42 12.8 4.501 5.094 121.9  9.5 15.3
8 1 38.15 −42 32.5 4.495 5.047 118.8  9.9 15.3
13 1 33.90 −42 46.9 4.493 4.999 115.5 10.3 15.3
18 1 29.82 −42 56.0 4.496 4.951 112.0 10.7 15.2

V1493 AQUILAE CCD magnitude estimates, unfiltered unless otherwise noted: Sept.

11.472 UT, 12.6 (S. Yoshida and K. Kadota, Ageo, Saitama, Japan); 15.793, V = 14.94 ± 0.01, B −V = +0.66 ± 0.07 (D. Hanžl, Brno, Czech Republic; cf. IAUC 7254); 26.7729, 13.83 (G. Masi, Ceccano, Italy); Oct. 1.761, V = 15.45 ± 0.05, B − V = +0.8 ± 0.10 (Hanžl).

1999 October 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green